# Ricci Curvature of Strong Product Graphs

### v1

@@ Abstract
We establish for the first time the explicit curvature formulas for the horizontal and vertical edges of the strong product of two regular graphs. We complement this result with showing that there does not exist a nice formula for the curvatures of diagonal edges, except for a special case in terms of curvatures of the factors. This gives the curvature formulas for all edges of the product of a complete graph and a regular graph. We also present an accessible and simpler proof of the curvature formulas for all the edges of the Cartesian product of two regular graphs, originally established by Lin, Lu, and Yau [2011].
@@ 1 Introduction
In the original paper, Lin et al. [12] studied the Cartesian product of two regular graphs. Cushing et al. [7] considered Ricci curvature of the strong product of two regular graphs. They proved lower curvature bounds for horizontal and vertical edges of $G\boxtimes H$ in terms of curvatures of the factors when $G$ and $H$ are regular graphs. The definitions will be deferred to the next section.
@@
The proof of Theorem 1.2 can be found in Section 3. Theorem 1.2 subsumes a special class of diagonal edges of $G\boxtimes H$. For general diagonal edge with endpoints $(x_{1},y_{1})$ and $(x_{2},y_{2})$, it is natural to wonder whether there exists an exact formula for $\kappa_{\rm{LLY}}((x_{1},y_{1}),(x_{2},y_{2}))$ which is purely in terms of $\kappa_{\rm{LLY}}(x_{1},x_{2})$, $\kappa_{\rm{LLY}}(y_{1},y_{2})$, $d_{G}$, and $d_{H}$. We give a simple example to explain that no such a formula exists at the end of Section 3.
@@ 2 Preliminaries
In the strong product of two graphs $G$ and $H$, denoted as $G\boxtimes H$, the vertex set is $V(G)\times V(H)$, and two distinct vertices $(x_{1},y_{1})$ and $(x_{2},y_{2})$ are adjacent if and only if $x_{2}\in N_{G}[x_{1}]$ and $y_{2}\in N_{H}[y_{1}]$. In other words, one of the following three conditions holds true: (i) $y_{1}y_{2}\in E(H)$ and $x_{1}=x_{2}$, (ii) $x_{1}x_{2}\in E(G)$ and $y_{1}=y_{2}$, or (iii) $x_{1}x_{2}\in E(G)$ and $y_{1}y_{2}\in E(H)$. The three types of edges are called, respectively, horizontal, vertical, and diagonal edges. In the Cartesian product of two graphs $G$ and $H$, denoted as $G\Box H$, the vertex set is $V(G)\times V(H)$, and $(x_{1},y_{1})$ and $(x_{2},y_{2})$ are adjacent if and only if the condition (i) or (ii) holds.
@@
Here $G*H$ refers to $G\boxtimes H$ or $G\Box H$, and the subscripts of OPT and MAX indicate on which graph we consider the value of the expression on the right.
@@ 3 Ricci curvature of strong product
The next two lemmas establish the relations between  $\text{OPT}_{G\boxtimes H}$ and $\text{OPT}_{H}$, and between $\text{MAX}_{H}$ and $\text{MAX}_{G\boxtimes H}$, respectively. In product graph, the subgraph induced by the vertex set $\{(x,y)\mid y\in H\}$ is called an $H$-fiber, and is denoted by ${}^{x}H$.
@@
Let $G$ and $H$ be two regular graphs with vertex degree $d_{G}$ and $d_{H}$, respectively. Let $x_{1},x_{2}\in V(G)$ with $N_{G}[x_{1}]=N_{G}[x_{2}]$, and $y_{1}y_{2}\in E(H)$. Then
@@
If $y_{1}y_{2}\in E(H)$ and $R_{y_{1}}(y_{1},y_{2})\neq\emptyset$. Since $x_{1}=x_{2}$ or $N_{G}[x_{1}]=N_{G}[x_{2}]$, it follows directly from Theorem 1.2. If $y_{1}=y_{2}$ or $R_{y_{1}}(y_{1},y_{2})=\emptyset$. By definition,
@@
Let $G=(V,E)$ be a locally finite graph, and $x,y\in V$ be two adjacent vertices with $d_{y}|d_{x}$. Then
@@
Recall that Theorem 1.2 subsumes a special class of diagonal edges of $G\boxtimes H$. We end this section by giving an example to show that for general diagonal edge with endpoints $(x_{1},y_{1})$ and $(x_{2},y_{2})$ in $G\boxtimes H$, $\kappa_{\rm{LLY}}((x_{1},y_{1}),(x_{2},y_{2}))$ cannot be uniquely determined by $\kappa_{\rm{LLY}}(x_{1},x_{2})$, $\kappa_{\rm{LLY}}(y_{1},y_{2})$, $d_{G}$, and $d_{H}$. As a consequence, no exact formula for $\kappa_{\rm{LLY}}((x_{1},y_{1}),(x_{2},y_{2}))$ purely in terms of $\kappa_{\rm{LLY}}(x_{1},x_{2})$, $\kappa_{\rm{LLY}}(y_{1},y_{2})$, $d_{G}$, and $d_{H}$ exists.
Let $G=C_{4}$, $H_{1}$ and $H_{2}$ be the 3-regular graphs with $y_{1}y_{2}$ being an edge in them, depicted in Figure 1. Let $x_{1}x_{2}$ be any one edge of $G$. By equation (3), $\kappa_{\rm{LLY}}^{H_{1}}(y_{1},y_{2})=\kappa_{\rm{LLY}}^{H_{2}}(y_{1},y_{2})=0$; however, $\kappa_{\rm{LLY}}^{G\boxtimes H_{1}}((x_{1},y_{1}),(x_{2},y_{2}))=-1/11$, $\kappa_{\rm{LLY}}^{G\boxtimes H_{2}}((x_{1},y_{1}),(x_{2},y_{2}))=0$ ($OPT_{G\boxtimes H_{1}}=13$, $OPT_{G\boxtimes H_{2}}=12$).
@@ 4 Ricci curvature of Cartesian product
Let $\phi_{1}\in\mathcal{O}_{(x_{1},y_{1})(x_{1},y_{2})}$, and $\phi_{2}\in\mathcal{O}_{y_{1}y_{2}}$. We claim that there exists some $\phi^{\prime}_{1}\in\mathcal{O}_{(x_{1},y_{1})(x_{1},y_{2})}$ such that for any $(x,y_{1})\in(N_{G}(x_{1})\times\{y_{1}\})$, $\phi^{\prime}_{1}((x,y_{1}))=(x,y_{2})$. To prove the claim, due to the optimality of $\phi^{\prime}_{1}$, it suffices to prove that there exists some $\phi^{\prime}_{1}\in\mathcal{O}_{(x_{1},y_{1})(x_{1},y_{2})}$ such that $\phi^{\prime}_{1}(N_{G}(x_{1})\times\{y_{1}\})=N_{G}(x_{1})\times\{y_{2}\}$. If $\phi_{1}$ satisfies this property, then let $\phi^{\prime}_{1}=\phi_{1}$ and we are done. Suppose that $\phi_{1}$ does not. Without loss of generality, assume that there exists only one vertex $(x^{\prime},y_{1})\in(N_{G}(x_{1})\times\{y_{1}\})$ such that $\phi_{1}((x^{\prime},y_{1}))\notin(N_{G}(x_{1})\times\{y_{2}\})$, i.e., $\phi_{1}((x^{\prime},y_{1}))\in(\{x_{1}\}\times R_{y_{2}}(y_{1},y_{2}))$. Let $\phi_{1}((x^{\prime},y_{1}))=(x_{1},y^{\prime})$. Since $|N_{G}(x_{1})\times\{y_{1}\}|=|N_{G}(x_{1})\times\{y_{2}\}|$. There must exist some $(x_{1},y^{\prime\prime})\in(\{x_{1}\}\times R_{y_{1}}(y_{1},y_{2}))$ such that $\phi_{1}((x_{1},y^{\prime\prime}))\in(N_{G}(x_{1})\times\{y_{2}\})$. Let $\phi_{1}((x_{1},y^{\prime\prime}))=(x^{\prime\prime},y_{2})$. Define a new assignment $\phi^{\prime}_{1}\in\mathcal{A}_{(x_{1},y_{1})(x_{1},y_{2})}$ as

### v2

@@ Abstract
We establish for the first time the explicit curvature formulas for the horizontal and vertical edges of the strong product of two regular graphs. We complement this result with showing that there does not exist an analogous formula for the curvatures of diagonal edges except for a special case, and providing a sharp lower bound for them in terms of the curvatures of the factors. This gives the curvature formulas for all the edges of the product of a complete graph and a regular graph. We also present an accessible and simpler proof of the curvature formulas for all the edges of the Cartesian product of two regular graphs, originally established by Lin, Lu, and Yau [2011].
@@ 1 Introduction
In the original paper, Lin et al. [12] studied the Cartesian product of two regular graphs. Cushing et al. [7] considered Ricci curvature of the strong product of two regular graphs. They proved lower curvature bounds for horizontal and vertical edges of $G\boxtimes H$ in terms of curvatures of the factors when $G$ and $H$ are regular graphs. In addition, Zhang and Fang [15] proved that the Lin–Lu–Yau curvature for strong product of two regular graphs can be (upper and lower) bounded by the curvatures of the factors. The definitions will be deferred to the next section.
@@
The proof of Theorem 1.2 can be found in Section 3. Theorem 1.2 subsumes a special class of diagonal edges of $G\boxtimes H$. For diagonal edge with endpoints $(x_{1},y_{1})$ and $(x_{2},y_{2})$ such that $N_{G}[x_{1}]\neq N_{G}[x_{2}]$ and $N_{H}[y_{1}]\neq N_{H}[y_{2}]$, it is natural to wonder whether there exists an exact formula for $\kappa_{\rm{LLY}}((x_{1},y_{1}),(x_{2},y_{2}))$ which is purely in terms of $\kappa_{\rm{LLY}}(x_{1},x_{2})$, $\kappa_{\rm{LLY}}(y_{1},y_{2})$, $d_{G}$, and $d_{H}$. We give a simple example to explain that no such a formula exists. Further we prove a sharp lower bound for $\kappa_{\rm{LLY}}((x_{1},y_{1}),(x_{2},y_{2}))$ in terms of  $\kappa_{\rm{LLY}}(x_{1},x_{2})$ and $\kappa_{\rm{LLY}}(y_{1},y_{2})$.
Theorem 1.3.
Let $G$ and $H$ be two regular graphs with vertex degrees $d_{G}$ and $d_{H}$, respectively. For any pair of adjacent vertices $x_{1}$ and $x_{2}$ of $G$ and any pair of adjacent vertices $y_{1}$ and $y_{2}$ of $H$, where $N_{G}[x_{1}]\neq N_{G}[x_{2}]$ and $N_{H}[y_{1}]\neq N_{H}[y_{2}]$,
$\kappa_{\rm{LLY}}((x_{1},y_{1}),(x_{2},y_{2}))\geq\frac{d_{H}(d_{G}+1)}{d_{G\boxtimes H}}\kappa_{\rm LLY}(y_{1},y_{2})+\frac{d_{G}(d_{H}+1)}{d_{G\boxtimes H}}\kappa_{\rm LLY}(x_{1},x_{2})-X,$
where
$X=1-\frac{|R_{x_{1}}(x_{1},x_{2})||R_{y_{1}}(y_{1},y_{2})|-1}{d_{G\boxtimes H}}.$
@@ 2 Preliminaries
In the strong product of two graphs $G$ and $H$, denoted as $G\boxtimes H$, the vertex set is $V(G)\times V(H)$, and two distinct vertices $(x_{1},y_{1})$ and $(x_{2},y_{2})$ are adjacent if and only if $x_{2}\in N_{G}[x_{1}]$ and $y_{2}\in N_{H}[y_{1}]$. In other words, one of the following three conditions holds true: (i) $y_{1}y_{2}\in E(H)$ and $x_{1}=x_{2}$, (ii) $x_{1}x_{2}\in E(G)$ and $y_{1}=y_{2}$, or (iii) $x_{1}x_{2}\in E(G)$ and $y_{1}y_{2}\in E(H)$. These conditions lead to three kinds of edges in product graph, which are called, respectively, horizontal, vertical, and diagonal edges. In the Cartesian product of two graphs $G$ and $H$, denoted as $G\Box H$, the vertex set is $V(G)\times V(H)$, and $(x_{1},y_{1})$ and $(x_{2},y_{2})$ are adjacent if and only if the condition (i) or (ii) holds.
@@
Here $G*H$ refers to $G\boxtimes H$ or $G\Box H$, and the subscripts of OPT and MAX indicate on which graph we consider the value of the expression on the right. Similarly, we can define $\text{OPT}_{G}$ and $\text{MAX}_{G}$.
@@ 3 Ricci curvature of strong product
The next two lemmas establish the relations between  $\text{OPT}_{G\boxtimes H}$ and $\text{OPT}_{H}$, and between $\text{MAX}_{H}$ and $\text{MAX}_{G\boxtimes H}$, respectively. In product graph, the subgraph induced by the vertex set $\{(x,y)\mid y\in V(H)\}$ is called an $H$-fiber, and is denoted by ${}^{x}H$.
@@
If $R_{y_{1}}(y_{1},y_{2})\neq\emptyset$. Since $N_{G}[x_{1}]=N_{G}[x_{2}]$, the result follows directly from Theorem 1.2.
If $R_{y_{1}}(y_{1},y_{2})=\emptyset$ or $y_{1}=y_{2}$. Since $R_{x_{1}}(x_{1},x_{2})=\emptyset$, by definition
@@
Let $G=(V,E)$ be a locally finite graph, and $x,y\in V$ with $xy\in E$ and $d_{y}|d_{x}$. Then
@@
In what follows, we give an example to show that in $G\boxtimes H$, for diagonal edge with endpoints $(x_{1},y_{1})$ and $(x_{2},y_{2})$ such that $N_{G}[x_{1}]\neq N_{G}[x_{2}]$ and $N_{H}[y_{1}]\neq N_{H}[y_{2}]$, $\kappa_{\rm{LLY}}((x_{1},y_{1}),(x_{2},y_{2}))$ cannot be uniquely determined by $\kappa_{\rm{LLY}}(x_{1},x_{2})$, $\kappa_{\rm{LLY}}(y_{1},y_{2})$, $d_{G}$, and $d_{H}$. As a consequence, no exact formula for $\kappa_{\rm{LLY}}((x_{1},y_{1}),(x_{2},y_{2}))$ purely in terms of $\kappa_{\rm{LLY}}(x_{1},x_{2})$, $\kappa_{\rm{LLY}}(y_{1},y_{2})$, $d_{G}$, and $d_{H}$ exists.
Let $G=C_{4}$, $H_{1}$ and $H_{2}$ be the 3-regular graphs with $y_{1}y_{2}$ being an edge, depicted in Figure 1. Let $x_{1}x_{2}$ be any one edge of $G$. By equation (3), it can be computed that $\kappa_{\rm{LLY}}^{H_{1}}(y_{1},y_{2})=\kappa_{\rm{LLY}}^{H_{2}}(y_{1},y_{2})=0$; however, $\kappa_{\rm{LLY}}^{G\boxtimes H_{1}}((x_{1},y_{1}),(x_{2},y_{2}))=-1/11$, $\kappa_{\rm{LLY}}^{G\boxtimes H_{2}}((x_{1},y_{1}),(x_{2},y_{2}))=0$ ($OPT_{G\boxtimes H_{1}}=13$, $OPT_{G\boxtimes H_{2}}=12$).
@@
Next we prove Theorem 1.3.
Proof of Theorem 1.3.
By definition of strong product, we have
$R_{(x_{1},y_{1})}((x_{1},y_{1}),(x_{2},y_{2}))$ $=(R_{x_{1}}(x_{1},x_{2})\times N_{H}[y_{1}])\cup(N_{G}[x_{1}]\times R_{y_{1}}(y_{1},y_{2})),$
$R_{(x_{2},y_{2})}((x_{1},y_{1}),(x_{2},y_{2}))$ $=(R_{x_{2}}(x_{1},x_{2})\times N_{H}[y_{2}])\cup(N_{G}[x_{2}]\times R_{y_{2}}(y_{1},y_{2})).$
Let $\phi_{1}\in\mathcal{O}_{x_{1}x_{2}}$, and $\phi_{2}\in\mathcal{O}_{y_{1}y_{2}}$. Define $\phi_{3}\in\mathcal{A}_{(x_{1},y_{1})(x_{2},y_{2})}$ as
$\phi_{3}(x,y)=\begin{cases}(x,\phi_{2}(y))&\text{if $(x,y)\in(N_{G}[x_{1}]\setminus R_{x_{1}}(x_{1},x_{2}))\times R_{y_{1}}(y_{1},y_{2})$,}\\ (\phi_{1}(x),y)&\text{if $(x,y)\in R_{x_{1}}(x_{1},x_{2})\times(N_{H}[y_{1}]\setminus R_{y_{1}}(y_{1},y_{2}))$,}\\ (\phi_{1}(x),\phi_{2}(y))&\text{if $(x,y)\in R_{x_{1}}(x_{1},x_{2})\times R_{y_{1}}(y_{1},y_{2})$.}\end{cases}$
By Proposition 3.2, it is easy to get
$\sum_{z\in R_{(x_{1},y_{1})}((x_{1},y_{1}),(x_{2},y_{2}))}d(z,\phi_{3}(z))$
$\leq\Big{(}(d_{G}+1)\sum_{z\in R_{x_{1}}(x_{1},x_{2})}d(z,\phi_{1}(z))+(d_{H}+1)\sum_{z\in R_{y_{1}}(y_{1},y_{2})}d(z,\phi_{2}(z))\Big{)}-|R_{x_{1}}(x_{1},x_{2})||R_{y_{1}}(y_{1},y_{2})|$
$=(d_{G}+1)\text{OPT}_{H}+(d_{H}+1)\text{OPT}_{G}-|R_{x_{1}}(x_{1},x_{2})||R_{y_{1}}(y_{1},y_{2})|,$
where the first inequality holds since for each $(x,y)\in R_{x_{1}}(x_{1},x_{2})\times R_{y_{1}}(y_{1},y_{2})$,
$d((x,y),\phi_{3}((x,y)))$ $=d((x,y),(\phi_{1}(x),\phi_{2}(y)))$
$=\max\left\{d(x,\phi_{1}(x)),d(y,\phi_{2}(y))\right\}$
$\leq d(x,\phi_{1}(x))+d(y,\phi_{2}(y))-1.$
Thus, we have
$\text{OPT}_{G\boxtimes H}$ $\leq\sum_{z\in R_{(x_{1},y_{1})}((x_{1},y_{1}),(x_{2},y_{2}))}d(z,\phi_{3}(z))$
$\leq(d_{G}+1)\text{OPT}_{H}+(d_{H}+1)\text{OPT}_{G}-|R_{x_{1}}(x_{1},x_{2})||R_{y_{1}}(y_{1},y_{2})|.$ (19)
Applying equation (3) on $G$ and $H$, we get
$\text{OPT}_{H}$ $=d_{H}-d_{H}\kappa_{\rm{LLY}}(y_{1},y_{2})+1,$ (20)
$\text{OPT}_{G}$ $=d_{G}-d_{G}\kappa_{\rm{LLY}}(x_{1},x_{2})+1.$
Combining equations (3), (20) and inequality (19) finishes the proof. ∎
Note that from the proof of Theorem 1.3, the lower bound can be attained if $\text{OPT}_{G}=|R_{x_{1}}(x_{1},x_{2})|$, and $\text{OPT}_{H}=|R_{y_{1}}(y_{1},y_{2})|$. For example, let $G=H=C_{4}$.
@@ 4 Ricci curvature of Cartesian product
Let $\phi_{1}\in\mathcal{O}_{(x_{1},y_{1})(x_{1},y_{2})}$, and $\phi_{2}\in\mathcal{O}_{y_{1}y_{2}}$. We claim that there exists some $\phi^{\prime}_{1}\in\mathcal{O}_{(x_{1},y_{1})(x_{1},y_{2})}$ such that for any $(x,y_{1})\in(N_{G}(x_{1})\times\{y_{1}\})$, $\phi^{\prime}_{1}((x,y_{1}))=(x,y_{2})$. To prove the claim, due to the optimality of $\phi^{\prime}_{1}$, it suffices to prove that there exists some $\phi^{\prime}_{1}\in\mathcal{O}_{(x_{1},y_{1})(x_{1},y_{2})}$ such that $\phi^{\prime}_{1}(N_{G}(x_{1})\times\{y_{1}\})=N_{G}(x_{1})\times\{y_{2}\}$. If $\phi_{1}$ satisfies this property, then let $\phi^{\prime}_{1}=\phi_{1}$ and we are done. Suppose that $\phi_{1}$ does not. Without loss of generality, assume that there exists only one vertex $(x^{\prime},y_{1})\in(N_{G}(x_{1})\times\{y_{1}\})$ such that $\phi_{1}((x^{\prime},y_{1}))\notin(N_{G}(x_{1})\times\{y_{2}\})$, i.e., $\phi_{1}((x^{\prime},y_{1}))\in(\{x_{1}\}\times R_{y_{2}}(y_{1},y_{2}))$. Let $\phi_{1}((x^{\prime},y_{1}))=(x_{1},y^{\prime})$. Since $|N_{G}(x_{1})\times\{y_{1}\}|=|N_{G}(x_{1})\times\{y_{2}\}|$. There must exist some $(x_{1},y^{\prime\prime})\in(\{x_{1}\}\times R_{y_{1}}(y_{1},y_{2}))$ such that $\phi_{1}((x_{1},y^{\prime\prime}))\in(N_{G}(x_{1})\times\{y_{2}\})$. Let $\phi_{1}((x_{1},y^{\prime\prime}))=(x^{\prime\prime},y_{2})$. Define $\phi^{\prime}_{1}\in\mathcal{A}_{(x_{1},y_{1})(x_{1},y_{2})}$ as